# TransKD: Transformer Knowledge Distillation for Efficient Semantic Segmentation

### v3

@@ IV-B Implementation Details
Teacher and student transformer models. For the main part of our semantic segmentation experiments, we investigate knowledge distillation for compressing a large pre-trained SegFormer B2 to a smaller SegFormer B0 model without pre-training (unless specified), which has a similar number of parameters as ResNets used in previous distillation works [12, 9, 10]. SegFormer B0 achieves a speed of $47.6$ Frames Per Second (FPS) when running at a resolution of $512{\times}1024$ and therefore suits well for real-time applications [26]. We have also examined our framework with PVTv2 [25] and LVT [34] to study the generality of our approach.
Training setups and hyperparameters. When evaluating the performance of the knowledge distillation frameworks, we resize the images to $512{\times}1024$ for Cityscapes [30], $512{\times}910$ for ACDC [31], and $512{\times}512$ for Pascal VOC2012 [33], while the original image scale of $480{\times}640$ is kept for the NYUv2 dataset [32]. We use one NVIDIA GTX 1080Ti GPU when experimenting with the aforementioned input sizes and four NVIDIA A100 GPUs for larger input sizes. The number of epochs is set to be $1000$ on Cityscapes, ACDC, and Pascal VOC2012 datasets, and $1500$ on the NYUv2 dataset. Following the structure of SegFormer [26], the height and width of the output and the target segmentation maps are all rescaled with a ratio of $1/4$ compared to the input image. The performance of the baseline SegFormer model has a slight decay, which is reasonable due to the smaller size of the input image. The models in our work are trained using the AdamW optimizer [76] with the learning rate of $6{\times}10^{-5}$ and adjusted by a polynomial learning rate decay scheduler with a factor of $1.0$ by default. We adopt a batch size of $8$ on VOC2012 and $2$ on other datasets. Besides, the number of channels of the feature maps in the CSF fusion module $C$ is set to $64$.
@@ V-A Quantitative results
Effectiveness for different architectures. We then verify whether the effectiveness of our method is consistent for different architectures. With this intent, we compare the TransKD-Base method against the Knowledge Review (KR) framework [10] for different segmentation backbones on the Cityscapes dataset. The results in Tab. IV indicate that our approach is not strictly tied to a concrete model, but is consistently beneficial for different segmentation backbones. Specifically, our TransKD-Base method achieves $+5.18\%$, $+2.18\%$, and $+2.71\%$ improvements over the feature-map-only KR method, while using the SegFormer-B0, the PVTv2-B0 [25], and the Lite Vision Transformer (LVT) [34] model as the student, respectively. For the PVTv2-B0, we use PVTv2-B2 [25] as the teacher model. In the setting of using the SegFormer-B2 as the teacher and the LVT as the student (which have rather dissimilar structures), our proposed TransKD framework also clearly improves the segmentation quality. When comparing to the original student segmenter without any distillation, the gains are more evident with $12.72\%$, $10.34\%$, and $11.05\%$ in mIoU.
These results show that our TransKD framework can unleash the potential of the efficient transformer models, yielding sufficient improvements of the compact models.
@@
Network #Params (M) GFLOPs mIoU (%)
Teacher (SegFormer-B2) 27.36 113.84 76.49
Student (SegFormer-B0) 3.72 13.67 55.86
+Knowledge Review 4.35 16.17 63.40
+TransKD-Base (ours) 4.56 16.47 68.58 (+5.18)
Teacher (PVTv2-B2) 29.1 83.14 77.26
Student (PVTv2-B0) 7.53 47.45 58.35
+Knowledge Review 8.15 49.94 66.51
+TransKD-Base (ours) 8.37 50.24 68.69 (+2.18)
Teacher (SegFormer-B2) 27.36 113.84 76.49
Student (LVT) 3.84 18.06 57.49
+Knowledge Review 4.47 20.63 65.83
+TransKD-Base (ours) 4.69 20.99 68.54 (+2.71)
@@ V-B Ablation Studies and Comparisons
Table XI: Accuracy analysis on different sizes of input images.
@@
Influence of input size and pre-training weights. According to the original work [26], using large input images significantly improves the segmentation results of SegFormer. Besides, ImageNet pre-training is a ubiquitous way for transformer-based methods to compensate for the lack of inductive bias and speed up convergence. In Tab. XI, we examine the influence of the input size and pre-trained weights by comparing the results of Knowledge Review [10] and TransKD-Base for different configurations.
We first look at the networks without weight initialization via pre-training, (i.e., without “+ImN” in Tab. XI). In this case, the large image scale facilitates the efficacy of the overall semantic segmentation model while compromising the gain of TransKD relative to Knowledge Review and the student. At the image size of $512{\times}1024$, the TransKD-distilled student raises the accuracy by $12.72\%$ compared to the original non-pretrained student. With the full size of $1024{\times}2048$, our TransKD-Base attains the performance of $71.59\%$ with a less, yet still significant accuracy increase of $9.73\%$ in mIoU.
Table XII: Performance comparison with state-of-the-art distillation methods over different student segmentation networks on the Cityscapes validation set [30]. * denotes that the efficient segmentation model uses pre-trained weights. “-” denotes that the corresponding information is missing in the respective paper.
@@
Comparison to state-of-the-art segmenters. In Tab. XII, we further compare TransKD-Base against state-of-the-art semantic segmentation networks and previous knowledge distillation methods. While heavier models like CCNet [78] and SETR [7] reach high mIoU scores, they are computationally intensive, which often disqualifies their deployment in mobile and real-time systems. Accurate segmenters like HANet [79] and SegFormer-B5 [26] require ${>}1000G$ FLOPs to yield precise high-resolution predictions. We further consider more efficient semantic segmenters. Compact models, such as CGNet [83] and Fast-SCNN [84], are very lightweight but come at a high accuracy cost, and the segmentation quality declines. Some models with moderate computation complexity like ERFNet [1] and BiSeNet [45] can attain higher accuracy but also comprise more model parameters.
Our previous experiments in Sec. V-A examined different knowledge distillation methods with the same student backbone. In this section, we conduct more comparisons to previously published methods with numbers from their respective papers for the enhanced student after distillation (Tab. XII). Most of the previous works have used PSPNet18 [27] as the student model. In the group of distillation methods and non-pretrained students, our TransKD-Base yields a top performance of $71.59\%$ in mIoU, which is even close to the results of some pre-trained students, e.g., MD [86] ($71.90\%$). In the group of distillation methods and pre-trained students, TransKD achieves the highest mIoU score despite being the most lightweight model with only $3.72M$ parameters and $54.70G$ FLOPs. Fig. 9 visualizes the comparison of representative efficient segmentation models and knowledge distillation methods. We clearly see that TransKD stands out in front of the past state-of-the-art models in both groups and strikes an excellent trade-off between efficiency and accuracy for semantic scene understanding.

### v4

@@ IV-B Implementation Details
Teacher and student transformer models. For the main part of our semantic segmentation experiments, we investigate knowledge distillation for compressing a large pre-trained SegFormer B2 to a smaller SegFormer B0 model without pre-training (unless specified), which has a similar number of parameters as ResNets used in previous distillation works [4, 1, 2]. SegFormer B0 [26] achieves a speed of $35.06$ Frames Per Second (FPS) when running at a resolution of $512{\times}1024$ on e NVIDIA GTX 1080Ti and is therefore well-suited for real-time applications. We have also examined our framework with PVTv2 [25] and LVT [34] to assess the generality of our approach.
Training setups and hyperparameters. When evaluating the performance of the knowledge distillation frameworks, we resize the images to $512{\times}1024$ for Cityscapes [30], $512{\times}910$ for ACDC [31], and $512{\times}512$ for Pascal VOC2012 [33], while maintaining the original image scale of $480{\times}640$ for the NYUv2 dataset [32]. We utilize one NVIDIA GTX 1080Ti GPU when experimenting with the aforementioned input sizes and four NVIDIA A100 GPUs for larger input sizes. The number of epochs is set to be $1000$ on Cityscapes, ACDC, and Pascal VOC2012 datasets, and $1500$ on the NYUv2 dataset. Following the structure of SegFormer [26], the height and width of the output and the target segmentation maps are all rescaled with a ratio of $1/4$ compared to the input image. The performance of the baseline SegFormer model has a slight decay, which is reasonable due to the smaller size of the input image. The models in our work are trained using the AdamW optimizer [76] with the learning rate of $6{\times}10^{-5}$, and the default epsilon of $1\times 10^{-8}$ and betas $(0.9,0.999)$. The learning rate is adjusted using a polynomial learning rate decay scheduler [77] with a default factor of $1.0$. The end learning rate is set to $0.0$, and the maximum number of decay steps is set at $1500$. We adopt a batch size of $8$ per GPU on VOC2012 and $2$ on other datasets. Additionally, the number of channels of the feature maps in the CSF fusion module $C$ is set to $64$.
@@ V-A Quantitative results
Effectiveness for different architectures. We then verify whether the effectiveness of our method is consistent for different architectures. With this intent, we compare the TransKD-Base method against the Knowledge Review (KR) framework [2] for different segmentation backbones on the Cityscapes dataset. The results in Tab. IV indicate that our approach is not strictly tied to a concrete model, but is consistently beneficial for different segmentation backbones. Specifically, our TransKD-Base method achieves $+5.18\%$, $+2.18\%$, and $+2.71\%$ improvements over the feature-map-only KR method, while using the SegFormer-B0, the PVTv2-B0 [25], and the Lite Vision Transformer (LVT) [34] model as the student, respectively. For the PVTv2-B0, we use PVTv2-B2 [25] as the teacher model. In the setting of using the SegFormer-B2 as the teacher and the LVT as the student (which have rather dissimilar structures), our proposed TransKD framework also clearly improves the segmentation quality. When comparing to the original student segmenter without any distillation, the gains are more evident with $12.72\%$, $10.34\%$, and $11.05\%$ in mIoU. These results show that our TransKD framework can unleash the potential of the efficient transformer models, yielding sufficient improvements of the compact models.
Furthermore, we assess the efficiency of various transformer architectures during both training and inference phases. During training, the complexity of teacher and student transformers embedded within knowledge distillation frameworks is measured in terms of the number of parameters and GFLOPs. Conversely, during the inference phase, the frame rate per second (FPS) is evaluated using pure transformers not equipped with knowledge distillation frameworks on the identical GPU, as the frameworks are primarily employed for knowledge transfer during training only. The student transformers utilized in this analysis are significantly smaller in size, ranging from $7.35$ to $3.86$ times smaller compared to the teacher models. Moreover, incorporating the distillation framework incurs only a minimal increase in the number of parameters (approximately $0.84$ million parameters) and computational cost (an additional $2.97$ GFLOPs), irrespective of the model type. Notably, the inference speed of the student models is enhanced by a factor of $2$ to $3.5$ times relative to that of the teacher models. Overall, student transformers offer faster inference with minimal additional cost and reduced sizes.
@@
Network #Params (M) GFLOPs FPS mIoU (%)
T: SegFormer-B2 27.36 113.84 10.17 76.49
S: SegFormer-B0 3.72 13.67 35.06 55.86
+KR 4.35 16.17 35.06 63.40
+TransKD-Base 4.56 16.47 35.06 68.58 (+5.18)
T: PVTv2-B2 29.1 83.14 9.91 77.26
S: PVTv2-B0 7.53 47.45 24.25 58.35
+KR 8.15 49.94 24.25 66.51
+TransKD-Base 8.37 50.24 24.25 68.69 (+2.18)
T: SegFormer-B2 27.36 113.84 10.17 76.49
S: LVT 3.84 18.06 20.42 57.49
+KR 4.47 20.63 20.42 65.83
+TransKD-Base 4.69 20.99 20.42 68.54 (+2.71)
@@ V-B Ablation Studies and Comparisons
Furthermore, we summarize the performances of TransKDs across various datasets in Tab. XI to analyze the effectiveness of the optimization modules, GL-Mixer and EA. We observe that the optimization modules consistently improve the performance of the basic version of TransKD by over $+0.5\%$. However, while TransKD enhanced by GL-Mixer underperforms the basic version on the ACDC dataset, TransKD-GL achieves state-of-the-art results in the snow condition, as demonstrated in Tab. V.
@@
Table XI: Accuracy analysis of optimization modules for patch embedding distillation.
Network Cityscapes ACDC NYUv2 VOC2012
Student 55.86 46.26 18.19 47.02
KR 63.40 54.90 22.80 61.64
TransKD-Base 68.58 58.56 23.61 65.81
TransKD-GL 68.87 58.13 24.20 66.13
TransKD-EA 68.98 59.09 23.89 66.32
Influence of input size and pre-training weights. According to the original work [26], using large input images significantly improves the segmentation results of SegFormer. Besides, ImageNet pre-training is a ubiquitous way for transformer-based methods to compensate for the lack of inductive bias and speed up convergence. In Tab. XII, we examine the influence of the input size and pre-trained weights by comparing the results of Knowledge Review [2] and TransKD-Base for different configurations.
We first look at the networks without weight initialization via pre-training, (i.e., without “+ImN” in Tab. XII). In this case, the large image scale facilitates the efficacy of the overall semantic segmentation model while compromising the gain of TransKD relative to Knowledge Review and the student. At the image size of $512{\times}1024$, the TransKD-distilled student raises the accuracy by $12.72\%$ compared to the original non-pretrained student. With the full size of $1024{\times}2048$, our TransKD-Base attains the performance of $71.59\%$ with a less, yet still significant accuracy increase of $9.73\%$ in mIoU.
Table XII: Accuracy analysis on different sizes of input images.
@@
Table XIII: Performance comparison with state-of-the-art distillation methods over different student segmentation networks on the Cityscapes validation set [30]. * denotes that the efficient segmentation model uses pre-trained weights. “-” denotes that the corresponding information is missing in the respective paper.
@@
Comparison to state-of-the-art segmenters. In Tab. XIII, we further compare TransKD-Base against state-of-the-art semantic segmentation networks and previous knowledge distillation methods. While heavier models like CCNet [79] and SETR [11] reach high mIoU scores, they are computationally intensive, which often disqualifies their deployment in mobile and real-time systems. Accurate segmenters like HANet [80] and SegFormer-B5 [26] require ${>}1000G$ FLOPs to yield precise high-resolution predictions. We further consider more efficient semantic segmenters. Compact models, such as CGNet [84] and Fast-SCNN [85], are very lightweight but come at a high accuracy cost, and the segmentation quality declines. Some models with moderate computation complexity like ERFNet [5] and BiSeNet [45] can attain higher accuracy but also comprise more model parameters.
Our previous experiments in Sec. V-A examined different knowledge distillation methods with the same student backbone. In this section, we conduct more comparisons to previously published methods with numbers from their respective papers for the enhanced student after distillation (Tab. XIII). Most of the previous works have used PSPNet18 [27] as the student model. In the group of distillation methods and non-pretrained students, our TransKD-Base yields a top performance of $71.59\%$ in mIoU, which is even close to the results of some pre-trained students, e.g., MD [87] ($71.90\%$). In the group of distillation methods and pre-trained students, TransKD achieves the highest mIoU score despite being the most lightweight model with only $3.72M$ parameters and $54.70G$ FLOPs. Fig. 9 visualizes the comparison of representative efficient segmentation models and knowledge distillation methods. We clearly see that TransKD stands out in front of the past state-of-the-art models in both groups and strikes an excellent trade-off between efficiency and accuracy for semantic scene understanding.
@@ V-C Qualitative Results
Limitation and future work. Although TransKDs enable non-pretrained student models to perform comparably to their pretrained counterparts, there remains a minor performance gap ($0.77\%$ on Cityscapes) between them. In the future, we anticipate that the knowledge sources, such as patch embeddings and feature maps, which have proven effective in our work, can be better leveraged to close this performance gap through more refined knowledge distillation framework designs. Furthermore, TransKDs have been exclusively implemented on pyramid transformers, which are primarily used in semantic segmentation. We believe that our TransKD framework could also be adapted to isotropic vision transformers.
@@ VI Conclusion
Appendix A Knowledge Distillation Loss Functions
Feature map loss function. Both patch embedding and feature map losses obtained from all four stages are weighted and added to the original cross-entropy loss. For feature map distillation, we compute the HCL loss [2] between the outputs of CSF and the teacher transformer block at each stage. The feature map obtained from CSF at each stage with shape $\left[n,c,h,w\right]$ is divided into $4$-level context information via adaptive average pooling. The height of the resulting multi-level feature maps at one stage is $\left[h,4,2,1\right]$, which denotes that the shapes of the four abstract feature maps are $\left[n,c,h,w\right]$, $\left[n,c,4,4\right]$, $\left[n,c,2,2\right]$, and $\left[n,c,1,1\right]$. At each stage, $L_{2}$ distances are utilized to distill knowledge between the context levels. The final HCL loss between the feature maps at the $m$-th stage is computed as:
$L_{fm}^{m}={\frac{1}{1+{\sum_{l}}{\frac{1}{2^{l}}}}}{\sum_{l=0}^{3}}{\frac{1}{2^{l}}}MSE(\mathbf{F}_{m,l}^{out},\mathbf{F}_{m,l}^{T}),$ (13)
where $l$ refers to the number of level at one stage. The purpose of placing the HCL loss after every stage is to encourage knowledge distillation at different levels of abstraction.
Appendix B More Performance Analyses
Analyses of per-class IoU improvement on various architectures. To further investigate the effectiveness of patch embedding distillation on various architectures, Fig. 12 illustrates per-class IoU improvements of our TransKD over Knowledge Review with all three teacher-student pairs, i.e. SegFormer B2-B0 [26], PVT B2-B0 [25], SegFormer B2-LVT [34]. Compared to Knowledge Review, TransKD helps the compact models to recognize less-frequent vehicle classes more accurately, such as truck, bus, and train.
Figure 12: Per-class IoU improvements of TransKD over Knowledge Review [2] with different teacher-student pairs.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Image Ground Truth Student TransKD-Base TransKD-GL TransKD-EA
Figure 13: Feature map visualizations. The feature maps at the last transformer encoder stage, i.e., the 4th stage, are visualized for analysis.
Image
Ground Truth
Teacher
Student
CD
KR
TransKD -Base
TransKD -GL
TransKD -EA
Normal Fog Night Rain Snow
Figure 14: Qualitative street scene semantic segmentation results under normal condition from the Cityscapes dataset [30] and adverse conditions from the ACDC dataset [31]. The performances of our knowledge distillation frameworks, TransKDs, are compared with Channel-wise Knowledge Distillation (CD) [1] and Knowledge Review (KR) [2].
Appendix C More Qualitative Analyses
Analyses of intermediate feature maps. We provide visual explanations of the network’s internal decision processes by visualizing the feature maps coming from the 4th stage of the model in Fig. 13, with representative Cityscapes validation set examples. Compared with the student without any TransKD components, TransKD-enhanced models present semantically recognizable responses. TransKD helps to pay more attention on the car and truck objects, as shown in the reddish regions on the feature maps, while maintaining clear boundaries with the background triggering much weaker feature map responses (blue and green regions), leading to better semantic segmentation outcome. Apart from that, TransKD-Base tends to focus on large objects, e.g., on the truck, while TransKD-GL shows more evident responses on cars which have relatively smaller size, which we attribute to the extraction of both local and global cues encouraged by this approach. Moreover, TransKD-EA better spotlights the objects which are far away from the camera, e.g., the car in the first row. These visualization results further illustrate the effectiveness of TransKD for learning semantic-aware intermediate representations.
Analyses of outdoor semantic segmentation. Here we compare the segmentation outputs of the most effective response-based distillation framework CD [1], feature-based distillation framework KR [2], and all three versions of TransKD. The outdoor semantic segmentation experiments are conducted on Cityscapes [30] (normal scenes) and ACDC dataset [31] (adverse scenes). According to the quantitative results, TransKD-GL shows the most satisfactory performance on the snowy weather, while TransKD-EA outperforms the all counterparts on normal and other adverse weathers. As shown in Fig. 14, the student SegFormer B0 without pre-training, and distilled with CD and KR, can roughly recognize the scenes, such as road, sidewalk, and vegetation, but the compact student transformer is still unable to segment the details, e.g., the traffic light and pedestrian in the distance and the occluded bus. On the contrary, TransKDs extracts both spatial and sequential relations through the combination of feature map and patch embedding distillation, thereby improving the consistency of the bus. Moreover, TransKDs can effectively distinguish the people, rider, and bicycle in adverse weathers, even though they are far away.
Analyses of indoor semantic segmentation. In our final study, we turn to the indoor semantic segmentation results and illustrate the segmentation outcomes on the NYUv2 dataset [32] in Fig. 15. Compared with the segmentation results produced by the student without any distillation, the segmentation boundaries of TransKD are much clearer, and most of the details of the indoor scenes are preserved, especially for the bedroom scene, e.g., on pillow. For the bathroom example, similar results can be found, where, for example, towel is better recognized by our approaches. Apart from the objects with relatively small size, objects with less texture cues can be completely segmented, e.g., the wall in the bedroom. These qualitative examples consistently confirm good generalization of TransKD across diverse environments (indoor and outdoor, different weather conditions, etc.) as well as the benefits of knowledge transfer from both patch embeddings and feature maps in transformer-to-transformer knowledge distillation.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Image Ground Truth Student
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
TransKD-Base TransKD-GL TransKD-EA
Figure 15: Qualitative indoor scene semantic segmentation results of our TransKD methods on the NYUv2 dataset [32].